# Scalable Oversight for Superhuman AI via Recursive Self-Critiquing

### v2

@@ Abstract
As AI capabilities increasingly surpass human proficiency in complex tasks, current alignment techniques including SFT and RLHF face fundamental challenges in ensuring reliable oversight. These methods rely on direct human assessment and become untenable when AI outputs exceed human cognitive thresholds. In response to this challenge, we explore two hypotheses: (1) critique of critique can be easier than critique itself, extending the widely-accepted observation that verification is easier than generation to the critique domain, as critique itself is a specialized form of generation; (2) this difficulty relationship is recursively held, suggesting that when direct evaluation is infeasible, performing high-order critiques (e.g., critique of critique of critique) offers a more tractable supervision pathway. To examine these hypotheses, we perform Human-Human, Human-AI, and AI-AI experiments across multiple tasks. Our results demonstrate encouraging evidence supporting these hypotheses and suggest that recursive self-critiquing is a promising direction for scalable oversight.
Machine Learning, ICML
footnotetext: The main idea for this work came from a late-night, insightful discussion between Jie Lou and Xing Yu , which was part of some truly wonderful days.footnotetext: This work was conducted during Xueru Wen and Xingyu Lu’s internship at Xiaohongshu.
@@ 1 Introduction
The provision of supervision signals is fundamental to AI alignment (Bowman et al., 2022). From the supervision signal acquisition perspective, tasks can be categorized as: (1) tasks with well-defined criteria, where ground truth can be deterministically obtained with low computational overhead, e.g., Go games and mathematical problems (Silver et al., 2017; Lightman et al., 2023); (2) tasks with subjectivity or complex evaluation frameworks, such as business strategy and product design jobs (Ouyang et al., 2022). The second type task is more prevalent in real-world applications and predominantly rely on human assessment, presenting a fundamental challenge for obtaining supervision signals.
Large language models achieve empirical success in alignment (Meta, 2024; Yang et al., 2024; DeepSeek-AI, 2024) with Supervised Fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) techniques. Specifically, SFT (Chung et al., 2022; Wei et al., 2022) finetunes models with human-annotated demonstrations, showing particular efficacy in tasks where humans can effectively showcase desired behaviors. RLHF (Christiano et al., 2023; Ouyang et al., 2022) leverages reinforcement learning with human preference reward models, extending supervision capability to more complex tasks where direct solution generation proves challenging.
@@
To systematically explore and verify these hypotheses, we first conduct Human-Human experiments, where human evaluate human outputs, examining the progression from response to critique and critique-of-critique ($C^{2}$). By comparing accuracy with similar effort, completion time, and confidence levels, we find that higher-order critique contributes to more effective evaluation than direct assessment. Furthermore, we demonstrate the recursive hypothesis by extending experimental settings to deeper critique chains, i.e., critique of critique of critique ($C^{3}$). Inspired by above human-human findings, we further investigate its applicability for supervising AI: when AI generate self-recursive critique, could humans provide effective oversight by evaluating these critique chains? To answer this question, we conduct Human-AI experiments, where human evaluate AI outputs, on tasks where AI outperforms average human. The result is promising and exhibited with models of varying capabilities. Finally, we examine whether AI could achieve effective oversight through self-recursive critiques, i.e., AI evaluate AI ouputs. Our results reveal that while contemporary AI systems demonstrate strong generation capabilities, they show notable limitations in higher-order criticism tasks.
@@
    Through comprehensive Human-Human, Human-AI, and AI-AI experiments, our findings provide valuable scalable oversight insights for surpervising advanced AI systems beyond human capabilities.
@@ 2.2 Baselines
Since higher-order critiques is based on all previous evaluation results, we need to consider computational cost for fair comparison. To validate that improvements stem from the recursive structure rather than increased computation, we compare higher-order critiques with majority voting of lower-order evaluations under similar effort. Specifically, let $\epsilon(\cdot)$ denote the computational overhead for each single evaluation. As presented in Figure 1, the total efforts $E(\cdot)$ for different order recursive critique $C^{1}$, $C^{2}$ and $C^{3}$ are defined as:
$E(C^{1})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1})\approx 3\epsilon(R),$ (5)
$E(C^{2})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1}_{1})+\epsilon(C^{1}_{2})+\epsilon(C^{2})$
$\approx 5\epsilon(R),$
$E(C^{3})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1}_{1})+\epsilon(C^{1}_{2})+\epsilon(C^{2}_{1})$
$+\epsilon(C^{2}_{2})+\epsilon(C^{3})\approx 7\epsilon(R).$
Then, we define majority voting at each level. For level $l$, given a set of $n$ evaluations, the majority voting result is:
@@
$C^{1}_{\text{naive}}(R_{1},R_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2}\}),$ (7)
$C^{2}_{\text{naive}}(C^{1}_{1},C^{1}_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2},A^{1}_{1},A^{1}_{2}\}),$
$C^{3}_{\text{naive}}(C^{2}_{1},C^{2}_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2},A^{1}_{1},A^{1}_{2},A^{2}_{1},A^{2}_{2}\}).$
@@ 3.1 Tasks
We select five representative tasks, which calls for diverse cognitive capabilities, and remain in moderate difficulty. All tasks include 64 multiple-choice questions.
@@
The questions are from China’s National Civil Service Examination (KAOGONG), the annual government recruitment assessment. These questions include logical reasoning, verbal comprehension, and quantitative analysis. We exclude general knowledge questions to focus on cognitive processing rather than factual recall.
@@
The questions are from the Civil Service Examination as well. These visual task assesses logical abilities through non-verbal reasoning without requiring extra domain knowledge or cultural context.
@@ 3.2 Setup
We recruit 32 participants with bachelor’s degrees, including 22 with STEM backgrounds and 10 with liberal arts backgrounds. Most participants passed CET-4 level English and achieved approximately 100 scores (out of 150) in high school math exam. Meawhile, most participants have full-time data annotation experience.
@@
We write standardized annotation guidelines for all tasks through descriptive instructions and examples, detailed in Appendix A. Tasks are distributed in data packages with specified submission deadlines, with random assignments across different critique levels to ensure all annotators participate in various stages. To maintain efficiency, we set a 20-minute time limit for each question at every critique stage. While time constraint exists, it is managed through package-level deadlines to allow flexible time allocation. Annotators complete a set number of tasks daily within their scheduled working hours. Feedback sessions are held to collect comments and suggestions for procedure or guideline improvement. We assign personnel for process management and annotation quality assurance.
@@ 3.3 Critique of Critique can be Easier than Critique
Table 3: Performance comparison across recursive criticism stages, with human accuracy subscripts showing difference from previous-stage AI accuracy. Results from two model scales (Qwen2.5-7B/72B) on mathematics and English tests, including accuracy, confidence (1-5), and completion time (minutes).
@@
3.4 Recursive Critique Remains Consistently Easier
We extend the recursive criticism to the third-order critique ($C^{3}$) on two representative tasks. As shown in Table 2, accuracy improves continuously at the $C^{3}$ level in both tasks, with CET-6 increasing from 60.94% at $C^{2}$ to 67.19%, and GAOKAO Math from 90.62% to 93.75%. More importantly, under comparable computational effort, majority voting shows similar improvements, reaching 67.19% for CET-6 and 93.75% for GAOKAO Math at $C^{3}$ level. Furthermore, naive voting achieves substantially lower performance than regular accuracy. These comparisons demonstrate that the effectiveness of recursive critique extends beyond mere computational scaling or consensus aggregation. Improving confidence scores and decreasing completion time further validate that recursive critique remains consistently easier.
4 Can Recursive Self-Critiquing Enable Human Oversight of AI?
In this section, we examine whether recursive critique enables effective human oversight when capabilities exceed human performance by asking human to evaluate AI-generated outputs and AI self-criticisms.
@@ 4.1 Tasks
We select tasks that AI demonstrates superior performance than our human annotators’ performance. The GAOKAO Math comprises the last two multiple-choice questions from high school mathematics examination (Zhang et al., 2023), while the TEM4 (Test for English Majors Grade Four) include reading comprehension questions. Mathematics questions demand advanced problem-solving skills, and TEM-4 questions require professional-level English proficiency. Both tasks are beyond most annotators’ abilities.
@@ 4.2 Setup
We employ the same annotators and evaluation metrics (accuracy, completion time, and confidence scores) as in Human-Human experiments. To generate AI responses, we utilize both Qwen-7B and Qwen-72B models (Qwen et al., 2025) to examine recursive criticism across different AI capability levels. We filter out questions on which models achieve either 0% or 100% accuracy, as these extremely easy or difficult tasks lead to uniform model outputs, making them unsuitable for validating recursive criticism. For each question, the AI model first outputs initial responses, then performs self-criticism recursively to certain orders ($C^{1}$, $C^{2}$). Human annotators only evaluate the final AI outputs. The annotation process follows the Human-Human procedure, except outputs are AI-generated.
@@ 4.3 Experimental Results
Table 3 indicates that humans’ response accuracy is lower than that of AI, showing that AI surpasses human in these tasks. However, for the following critique, humans consistently achieve higher accuracy than AI’s previous outputs. For example, with Qwen2.5-7B on GAOKAO Math, human accuracy reach 53.12% at the critique stage (7.03% higher than AI’s initial 46.09%), and further increases to 56.25% at $C^{2}$ (8.59% above AI’s critique stage). This finding suggests that recursive criticism enables human supervision even when AI outperforms human for generation.
@@
5 Can AI Perform Recursive Self-Critiquing?
To investigate the effectiveness of the Recursive Self-Critiquing protocol on existing AI models, we conduct experiments on a wide range of LLMs.

### v3

@@ Abstract
As AI capabilities increasingly surpass human proficiency in complex tasks, current alignment techniques including SFT and RLHF face fundamental challenges in ensuring reliable oversight. These methods rely on direct human assessment and become untenable when AI outputs exceed human cognitive thresholds. In response to this challenge, we explore two hypotheses: (1) Critique of critique can be easier than critique itself, extending the widely-accepted observation that verification is easier than generation to the critique domain, as critique itself is a specialized form of generation; (2) This difficulty relationship is recursively held, suggesting that when direct evaluation is infeasible, performing high-order critiques (e.g., critique of critique of critique) offers a more tractable supervision pathway. We further conduct Human-AI and AI-AI experiments to investigate the potential of utilizing recursive self-critiquing for AI supervision. Our results highlight recursive critique as a promising approach for scalable AI oversight.
@@ 1 Introduction
The provision of supervision signals is fundamental to AI alignment (Bowman et al., 2022). From the supervision signal acquisition perspective, tasks can be categorized as: (1) tasks with well-defined criteria, where ground truth can be deterministically obtained with low computational overhead, e.g., Go games and mathematical problems (Silver et al., 2017; Lightman et al., 2023); (2) tasks with subjectivity or complex evaluation frameworks, such as business strategy and product design jobs (Ouyang et al., 2022). Tasks of the second type are more prevalent in real-world applications and predominantly rely on human assessment, presenting a fundamental challenge.
Large language models achieve empirical success in alignment (Meta, 2024; Yang et al., 2024; DeepSeek-AI, 2024) through Supervised Fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) techniques. Specifically, SFT (Chung et al., 2022; Wei et al., 2022) finetunes models with human-annotated demonstrations, showing particular efficacy in tasks where humans can effectively showcase desired behaviors. RLHF (Christiano et al., 2023; Ouyang et al., 2022) leverages reinforcement learning with human preference reward models, extending supervision capability to more complex tasks where direct solution generation proves challenging.
@@
To systematically verify and explore these hypotheses, we first conduct Human-Human experiments, where human evaluate human outputs, examining the progression from response to critique and critique-of-critique ($C^{2}$). By comparing accuracy with similar effort, completion time, and confidence levels, we find that higher-order critique contributes to more effective evaluation than direct assessment. Furthermore, we demonstrate the recursive hypothesis by extending experimental settings to deeper critique chains, i.e., critique of critique of critique ($C^{3}$). Inspired by above human-human findings, we further investigate its applicability for supervising AI: when AI generate self-recursive critique, could humans provide effective oversight by evaluating these critique chains? To answer this question, we conduct Human-AI experiments, where human evaluate AI outputs, on tasks where AI outperforms average human. The result is promising and exhibited with models of varying capabilities. Finally, we examine whether AI could achieve effective oversight through recursive self-critiques in AI-AI experiments across models of different capabilities. Our results demonstrate that recursive self-critiquing shows effectiveness in weak-to-strong scenarios, while the optimal critique strategy depends on relative model strengths between the supervised and critic models.
@@
    Through comprehensive Human-Human, Human-AI, and AI-AI experiments, we demonstrate the potential of recursive self-critiquing as a scalable oversight method, providing new valuable insights for supervising advanced AI systems beyond human capabilities.
footnotetext: The main idea for this work came from a late-night, insightful discussion between Jie Lou and Xing Yu, which was part of some truly wonderful days.
@@ 2.2 Baselines
Since higher-order critiques are based on all previous evaluation results, we need to consider computational cost for fair comparison. To validate that improvements stem from the recursive structure rather than increased computation, we compare higher-order critiques with majority voting of lower-order evaluations under similar effort. Specifically, let $\epsilon(\cdot)$ denote the computational overhead for each single evaluation. As presented in Figure 1, the total efforts $E(\cdot)$ for different order recursive critique $C^{1}$, $C^{2}$ and $C^{3}$ are defined as:
$E(C^{1})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1})\approx 3\epsilon(R)$ (5)
$E(C^{2})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1}_{1})+\epsilon(C^{1}_{2})+\epsilon(C^{2})\approx 5\epsilon(R)$
$E(C^{3})$ $=\epsilon(R_{1})+\epsilon(R_{2})+\epsilon(C^{1}_{1})+\epsilon(C^{1}_{2})+\epsilon(C^{2}_{1})+\epsilon(C^{2}_{2})+\epsilon(C^{3})\approx 7\epsilon(R)$
Then, we define majority voting. For level $l$, given a set of $n$ evaluations, the majority voting result is:
@@
$C^{1}_{\text{naive}}(R_{1},R_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2}\})$ (7)
$C^{2}_{\text{naive}}(C^{1}_{1},C^{1}_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2},A^{1}_{1},A^{1}_{2}\})$
$C^{3}_{\text{naive}}(C^{2}_{1},C^{2}_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2},A^{1}_{1},A^{1}_{2},A^{2}_{1},A^{2}_{2}\})$
@@ 3.1 Tasks
We select five representative tasks, which call for diverse cognitive capabilities, and remain in moderate difficulty. All tasks include 64 multiple-choice questions.
@@
The questions are from China’s National Civil Service Exam, the annual government recruitment test. These questions cover logical reasoning, language understanding, and number analysis. We exclude knowledge questions to focus on cognitive skills rather than factual recall.
@@
These visual tasks from the Civil Service Examination assess logical abilities through non-verbal reasoning without requiring extra knowledge or cultural context.
@@ 3.2 Setup
We recruit 32 participants with bachelor’s degrees, including 22 with STEM backgrounds and 10 with liberal arts backgrounds. Most participants passed CET-4 level English and achieved a score of approximately 100 (out of 150) in high school math exam. Meawhile, most participants have full-time data annotation experience.
@@
We write standardized guidelines for all tasks through descriptive instructions and examples, detailed in Appendix A. Tasks are distributed in data packages with specified submission deadlines, with random assignments across different critique levels to ensure all annotators participate in all stages. To maintain efficiency, we set a 20-minute time limit for each question at every stage. While the time constraint exists, it is managed through package-level deadlines to allow flexible time allocation. Annotators complete a set number of tasks daily within their scheduled working hours. Feedback sessions are held to collect comments and suggestions for procedure or guideline improvement. We assign personnel for process management and annotation quality assurance.
@@ 3.3 Critique of Critique can be Easier than Critique
3.4 Recursive Critique Remains Consistently Easier
We extend the recursive criticism to the third-order critique ($C^{3}$) on two representative tasks. As shown in Table 2, accuracy improves continuously at the $C^{3}$ level in both tasks, with CET-6 increasing from 60.94% at $C^{2}$ to 67.19%, and GAOKAO Math from 90.62% to 93.75%. More importantly, under comparable computational effort, majority voting shows similar improvements, reaching 67.19% for CET-6 and 93.75% for GAOKAO Math at $C^{3}$ level. Furthermore, naive voting achieves substantially lower performance than regular accuracy. Meanwhile, confidence scores improve as completion time decreases. These results demonstrate that recursive critique remains consistently easier and extends beyond mere computational scaling or consensus aggregation.
4 Can Recursive Self-Critiquing Enable Human Oversight of AI?
In this section, we further conduct Human-AI experiments to examine whether recursive critique enables effective human oversight when capabilities exceed human performance.
Table 3: Performance comparison across recursive criticism stages, with human accuracy subscripts showing difference from previous-stage AI accuracy. Results from Qwen2.5-7B/72B-Instruct on mathematics and English tests, including accuracy, confidence (1-5), and completion time (minutes).
@@ 4.1 Tasks
We select tasks where AI demonstrates superior performance compared to our human annotators’ capabilities. Specifically, the GAOKAO Math comprises the last two multiple-choice questions from the high school mathematics examination (Zhang et al., 2023), while the TEM4 (Test for English Majors Grade Four) includes reading comprehension questions. Importantly, these mathematics questions demand advanced problem-solving skills, and similarly, TEM-4 questions require professional-level English proficiency. As a result, both tasks are beyond most annotators’ abilities.
@@ 4.2 Setup
We employ the same annotators and evaluation metrics as in Human-Human experiments. To generate AI responses, we utilize both Qwen-7B-Instruct and Qwen-72B-Instruct models (Qwen et al., 2025) to examine recursive criticism across different AI capability levels. We filter out questions on which models achieve either 0% or 100% accuracy, as these extremely easy or difficult tasks lead to uniform model outputs, making them unsuitable for validating recursive criticism. For each question, the AI model first outputs initial responses, then performs self-criticism recursively to certain orders ($C^{1}$, $C^{2}$). Human annotators only evaluate the final AI outputs. The annotation process follows the Human-Human procedure, except outputs are AI-generated.
@@ 4.3 Experimental Results
Table 3 indicates that humans’ response accuracy is lower than that of AI, showing that AI surpasses humans in these tasks. However, for the following critique, humans consistently achieve higher accuracy than AI’s previous outputs. For example, with Qwen2.5-7B on GAOKAO Math, human accuracy reaches 53.12% at the critique stage (7.03% higher than AI’s initial 46.09%), and further increases to 56.25% at $C^{2}$ (8.59% above AI’s critique stage). This finding suggests that recursive criticism enables human supervision even when AI outperforms humans for generation.
@@
5 Can Recursive Self-Critiquing Achieve Better AI Supervision?
In this section, we conduct AI-AI experiments to explore the potential of recursive self-critiquing for achieving better AI supervision under weak-to-strong, strong-to-weak, and self-supervised settings.